Converting the energy of one food calorie into mechanical energy, how high could you lift a 196-lbs person?
Answer #1

E = 1cal = 4.12 J, m = 196 lbs =88.9 kg

From conservation of energy

E = U+K

E = mgh +0

4.2 = (88.9*9.8)h

h = 0.0048 m